OpenStudy (anonymous):

(x-2)^(2/3)+(x-2)^(1/3)-12=0

myininaya (myininaya):

let u=(x-2)^(1/3) then u^2=(x-2)^(2/3) so we have u^2+u-12=0 (u+4)(u-3)=0 u=-4, u=3 but remember u=(x-2)^(1/3) so solve (x-2)^(1/3)=-4 and (x-2)^(1/3)=3 for x